OpenStudy (anonymous):

The range of a set of data is 75. If the lowest value in the data set is 5, what is the highest value? 15 25 80 375

OpenStudy (nathan917):

75 = h - l 75 = h - 5 Answer: h = 80

OpenStudy (texaschic101):

The range is the difference between the lowest number and the highest number. So if the lowest number is 5, then add 75 to it to get the highest number. 5 + 75 = 80 Voila !! easy :)